1. Find the domain of the expression x/(x^2 - 9).
x^2 - 9 = 0
x^2 = 9
sqrtx^2} = sqrt{9}
x = -3, and x = 3.
Domain = {x | x cannot be -3 and 3}.

2. Find the domain of the expression
(-9x^2 - x + 1)/(x^3 + x).
x^3 + x =
x(x^2 + 1) = 0
Setting x^2 + 1 to zero and solving for x leads to complex roots. I will disregard complex roots for now. The only value that x cannot be is 0.
Domain = {x | x cannot be 0}
What do y I say?

Answer by math_tutor2020(3817)


Problem 1.

x^2-9 = 0 had been correctly solved to get x = -3 and x = 3.
And you are correct that x cannot equal either of these values; otherwise, we get a division by zero error.

Another way to solve is:
x^2-9 = 0
x^2-3^2 = 0
(x-3)(x+3) = 0 .... difference of squares rule
x-3 = 0 or x+3 = 0
x = 3 or x = -3

The domain in set-builder notation is

Or you can slightly condense things to get

The portion translates to "x is in the set of real numbers" i.e. "x is a real number".


The interval notation would look like this

This is basically us starting with the entire number line and then poking two holes at x = -3 and x = 3 on the number line.
Each "U" refers to a set union joining or gluing together the disjoint intervals.
Curved parenthesis exclude each endpoint.

--------------------------------------------------------------------------

Problem 2.

You have the correct scratch work. I'll assume your second step should have read x^3+x=0.


The domain in set-builder notation is

In words we can say "x is any nonzero real number".


The domain as interval notation would be
It's the same idea as earlier, except we only have one hole this time.